OpenStudy (ashking1):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 4(x + 7)2(x - 7)3

OpenStudy (ashking1):

these are the choices 4, multiplicity 1; -7, multiplicity 3; 7, multiplicity 3 -7, multiplicity 3; 7, multiplicity 2 4, multiplicity 1; 7, multiplicity 1; -7, multiplicity 1 -7, multiplicity 2; 7, multiplicity 3 can someone help me please

OpenStudy (amoodarya):

do you mean ?\[f(x) = 4(x + 7)^2(x - 7)^3\]

OpenStudy (phi):

you should put in ^ to show exponents. I assume this is f(x) = 4(x + 7)^2(x - 7)^3 \[ f(x) = 4(x + 7)^2(x - 7)^3 \]

OpenStudy (ashking1):

yes

OpenStudy (phi):

what happens if you erase the x and put in 7 in its place in (x-7) what number do you get ?

OpenStudy (ashking1):

0

OpenStudy (ashking1):

3 times

OpenStudy (phi):

yes. they use the word "multiplicity" so look for an answer that says 7 multiplicity 3

OpenStudy (ashking1):

-7, multiplicity 3; 7, multiplicity 2 so would this be the answer

OpenStudy (phi):

it is easy to get confused. 4(x + 7)^2(x - 7)^3 we did (x-7) . what x makes that zero?

OpenStudy (ashking1):

what do you mean by that like its the second x the one that says in (x-7) and if x is 7 it will equal 0 right ?

OpenStudy (phi):

(x-7) is zero when x=7 and multiplicity 3 because we have (x-7)^3

OpenStudy (phi):

you picked *** 7, multiplicity 2 so would this be the answer***

OpenStudy (ashking1):

no

OpenStudy (ashking1):

it would be -7, multiplicity 2; 7, multiplicity 3

OpenStudy (phi):

yes

OpenStudy (ashking1):

thanks btw can you help me with one more

OpenStudy (phi):

ok

OpenStudy (ashking1):

i need help understanding either the triangle or the theorem to answer this question Expand the following using either the Binomial Theorem or Pascal’s Triangle. You must show your work for credit. (x - 5)5

OpenStudy (ashking1):

like you dont have to tell me the answer more like help me understand how to get there

OpenStudy (phi):

please use the ^ otherwise it looks like (x-5) times 5 you mean \( (x-5)^5 \) first, write it like this: \( (x + (-5) )^5 \)

OpenStudy (ashking1):

sorry ok

OpenStudy (phi):

next, we start with this pattern \[ (a+b)^5 = a b + a b + a b + ...\] make the exponent of the a in the first term 5, and the exponent on b 0 (the exponents will *always* add up to 5) the next a will have an exponent of 4 (and the b will have 1) can you write out the a b pattern for (a+b)^5 ?

OpenStudy (ashking1):

so it will be (5+0)^5= (4)(1).....

OpenStudy (phi):

no. I am trying to say that \[ (a+b)^5 = a^5 b^0 + a^4 b^1 + ... \] (there are also numbers we have to put in, but first we need to learn how to write down this pattern) can you finish writing out the full pattern ?

OpenStudy (ashking1):

would the next one be like a^3b^2

OpenStudy (phi):

yes keep going

OpenStudy (ashking1):

a^2b^2

OpenStudy (phi):

you mean a^2 b^3 (remember the exponents add up to 5) or the a goes down, the one on b goes up

OpenStudy (ashking1):

ya sorry thats what i ment

OpenStudy (phi):

keep going, we need all the terms

OpenStudy (ashking1):

then a^1B^4

OpenStudy (ashking1):

a^0b^5

OpenStudy (phi):

yes. so we have (so far) \[ (a+b)^5 = a^5 b^0 + a^4 b^1 + a^3b^2+a^2b^3 + a^1 b^4 +a^0 b^5\] now we have to put in coefficients (numbers) in front of each term. we use Pascal's triangle to find these numbers.

OpenStudy (phi):

start by writing this pattern 

OpenStudy (phi):

next we add a new row by adding 1 on each side, and adding the two numbers in the row above it. Like this